OpenStudy (konradzuse):

Find the area of the interior of r = 12sin(theta)

OpenStudy (lgbasallote):

sounds like an integral haha

OpenStudy (konradzuse):

:O? :) I like integrals.

OpenStudy (lgbasallote):

idk...i just said sounds like...usually when i see area i think of integrals =))

OpenStudy (lgbasallote):

i dont know about integrals of polar coordinates though :C

OpenStudy (anonymous):

first find the limits for theta solve 12sin(theta)=0

OpenStudy (lgbasallote):

so it is an integral @sami-21 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@lgbasallote is is just to find the area of circle whose radius is 6

OpenStudy (konradzuse):

:O hmm :)

OpenStudy (anonymous):

here is helpful graph.

OpenStudy (konradzuse):

;) yeah I saw wolfram. So woulod it be arctan(12) = theta?

OpenStudy (anonymous):

arcsin(0) :P

OpenStudy (konradzuse):

12 disappears?

OpenStudy (konradzuse):

I figured it swapped, but wasn't sure where the 12 goes.

OpenStudy (konradzuse):

so theta = 0 anyways....

OpenStudy (anonymous):

yes 0=12sin(theta) sin(theta)=0/12 theta=arcsin(0) !

OpenStudy (konradzuse):

Ah okay I thought about 0/12 too :P.

OpenStudy (konradzuse):

theta = 0.

OpenStudy (anonymous):

and where else sin is zero in the range of 0..2pi :P

OpenStudy (konradzuse):

pi :)

OpenStudy (anonymous):

yes !!!!

OpenStudy (konradzuse):

:D

OpenStudy (konradzuse):

so theta = pi?

OpenStudy (anonymous):

you know the integral for area in polar coordinates is. \[\Large Area=\int\limits_{\theta_{1}}^{\theta_{2}}\frac{1}{2}r^2d \theta\] so here r=12sin(theta) and you have found the limits for theta. so integral becomes \[\Large Area=\frac{1}{2}\int\limits_{0}^{\pi} (12\sin(\theta))^2d \theta\]

OpenStudy (konradzuse):

so we should change it to 144/2 = 72 int sin^2(theta) d theta

OpenStudy (konradzuse):

so for sin^2 should we change it using the double angel/half angel formulas?

OpenStudy (anonymous):

yes !!

OpenStudy (anonymous):

This wolfram thing has ruined the student to do integration by themselves :P Even simple integral cannot be solved by students themselves who are addicted to wolfram :(. When I was studying integrals there was no wolfram thing :P

OpenStudy (konradzuse):

:)

OpenStudy (konradzuse):

sorry just got back.

OpenStudy (konradzuse):

wolfram is good to check your answers or to ask basic questions.

OpenStudy (anonymous):

dit you get it now ?

OpenStudy (konradzuse):

yessir.. It asks me to graph it tho so Idk what to do with that...

OpenStudy (konradzuse):

"Graph and then find the area r = 12sin(theta)"

OpenStudy (konradzuse):

It gives me a circle from -6 to 6 for both x,y...

OpenStudy (konradzuse):

which is what the circle is.....

OpenStudy (konradzuse):

or should be from 0 to 12 for y.

OpenStudy (anonymous):

should be from zero to 12 on y. because sin(90)=1 so r=12 !!

OpenStudy (konradzuse):

let me see if I can post it...

OpenStudy (konradzuse):

@sami-21

OpenStudy (anonymous):

what you want me to do ? to check the graph ?

OpenStudy (konradzuse):

im not sure how to graph it... it should be 0-12

OpenStudy (konradzuse):

also I gpt 36x - 18sin(2x)

OpenStudy (anonymous):

yes it should be just try o,pi/2,pi,3pi/2 ,2pi in the given equation r=12sin(theta) and plot the values of r yo get from above angles . then connect those point.

OpenStudy (konradzuse):

should I do the formula sin(2u) = 2sin u cos u?

OpenStudy (konradzuse):

or can I leave it as that?

OpenStudy (konradzuse):

it should come out the same with the limits...

OpenStudy (anonymous):

nooooooooooooooooooooooooooooooooooooooooooo :P it is sin^2(x) not sin(2x) use \[\Large \sin^2(x)=\frac{(1-\cos(2x)}{2}\]

OpenStudy (konradzuse):

yeah that's up top... But I end up with 36x-18sin(2x)

OpenStudy (konradzuse):

I ended up geting 36pi.

OpenStudy (anonymous):

good that's correct !!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (konradzuse):

the reason I ask is when I checked with the evil wolfram it said it switch it over... http://www.wolframalpha.com/input/?i=integral+1%2F2+%2812sin%28x%29%29%5E2

OpenStudy (anonymous):

Just to check your answer ! I told you it is to find the area of circle whose radius is 6 you know formula for area is \[\Large Area=\pi r^2\] So \[\Large Area= \pi (6(^2)=36 \pi \]

OpenStudy (konradzuse):

oh cooool!

OpenStudy (anonymous):

so it is ok now :)

OpenStudy (konradzuse):

now to graph this stupid thing :(

OpenStudy (konradzuse):

so you said to plug in 0 pi etc for theta and plot huh....

OpenStudy (konradzuse):

:O?

OpenStudy (anonymous):

it is the equation of circle centered on y axis with radius r=6 the standard equation of circle centered at y axis with radius a is given by. \[\Large r=2asin(\theta)\]

OpenStudy (konradzuse):

hmmm...

OpenStudy (anonymous):

this means it will not comes down below y axis. study about polar graphs in your text.

OpenStudy (konradzuse):

so how do I graph this bs grrr!! :(

OpenStudy (konradzuse):

it should be all the wa to the top, do I stop at 6 like wtf do i DPOOO :(

OpenStudy (konradzuse):

maybe I just weont graph it LOL